In a shuffleboard game, the puck slides a total of 12 m before coming to rest. If the coefficient of kinetic friction between the puck and the horizontal board is 0.28, what was the initial speed of the puck?
1 : 8.1 m/s
2 : 29.0 m/s
3 : 6.5 m/s
4 : 7.3 m/s
1
Expert's answer
2020-05-28T12:01:15-0400

The acceleration of the puck

"a=\frac{v^2}{2d}=
\frac{F}{m}\\
\frac{v^2}{2d}=\mu g"

Hence


"v=\sqrt{2\mu g d}=\sqrt{2\times 0.28\times 9.8\times 12}=8.1\:\rm m/s"